OpenStudy (anonymous):

If the range of the function f(x)= 5x − 3 is {-13, -8, -3, 12, 17}, what is its domain? I am confused with the 5x-3, can someone help?

OpenStudy (anonymous):

@iCake @ICanAwnserQuestions

OpenStudy (anonymous):

i do not know.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The range is the set of possible outputs. So it's saying that the set of possible outputs are -13, -8, -3, 12, 17

OpenStudy (anonymous):

yeah i get that what i dont get is the 5x-3

jimthompson5910 (jim_thompson5910):

it wants the corresponding inputs for each output so say the output is y = -13 that means y = 5x - 3 -13 = 5x - 3 -13+3 = 5x -10 = 5x 5x = -10 x = -10/5 x = -2 So if the output is y = -13, then the input is x = -2

jimthompson5910 (jim_thompson5910):

you will do this for the other outputs to get the domain

OpenStudy (anonymous):

oh now i get it thanks!!! :D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what did you get? i get -2,-1,0,1,2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

I don't agree with 1 or 2 but the "-2,-1,0" part is correct

OpenStudy (anonymous):

okay let me try it again

OpenStudy (anonymous):

i dont get it what i have to do with the y=-13 and x-2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you see how I plugged in y = -13 and solved to get x = -2 ?

OpenStudy (anonymous):

yeah but i dont undersand it that much

jimthompson5910 (jim_thompson5910):

well the idea is that if you plug in x =-2, you'll get y = -13 out they want you to think in reverse. If y = -8 is the output, then what must the input be?

OpenStudy (anonymous):

-1 ?

OpenStudy (anonymous):

i mean -4

jimthompson5910 (jim_thompson5910):

yep because y = 5x-3 y = 5*(-1) - 3 ... replace x with -1 y = -5 - 3 y = -8 so this shows how if x = -1 is plugged in, then y = -8 pops out

jimthompson5910 (jim_thompson5910):

I'll be right back

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

I'm back

jimthompson5910 (jim_thompson5910):

So do you see how plugging in x = -1 gives y = -8?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

and if y = -3, then what must x be?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

let's check y = 5x-3 y = 5*2 - 3 ... replace x with 2 y = 10-3 y = 7 so if x = 2, then y = 7 but we want y = -3

OpenStudy (anonymous):

i get it! thank you! I have 4 questions more can you help me? its for a home test

jimthompson5910 (jim_thompson5910):

I can't help with tests

OpenStudy (anonymous):

its practically a homework and its the last 4 questions that i have trouble with :/

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Ok I'll help with 2 of them if it's a hw. Which 2 are giving you the most trouble?

OpenStudy (anonymous):

widget wonder produce widgets. They have found that the cost,c(x), of making x widgets is a quadratic function in terms of x. The company also discovered that it cost $16 to produce 2 widgets, $18 to produce 4 widgets, and $48 to produce 10 widgets. Find the total cost of producing 6 widgets.

OpenStudy (anonymous):

i am so bad at that kind of problems like the paraghrap @jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok let me look through it

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

do you have a TI calculator (like a TI83 or TI84) ?

OpenStudy (anonymous):

no :/ but cellphone work? or a scientific calculator?

jimthompson5910 (jim_thompson5910):

the cellphone won't work because we need to do quadratic interpolation

jimthompson5910 (jim_thompson5910):

let me find a calculator

OpenStudy (anonymous):

okay thanks for your effort and your time ;) :D

jimthompson5910 (jim_thompson5910):

ok I found this page http://www.solvemymath.com/online_math_calculator/interpolation.php you will type in [2,16], [4,18], [10,48] and hit calculate. Tell me what comes up

jimthompson5910 (jim_thompson5910):

I got that as well

jimthompson5910 (jim_thompson5910):

now you plug x = 6 into that function to get your final answer

OpenStudy (anonymous):

thas what i dont get because it a big division

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = \frac{x^2-4x+36}{2}\] \[\Large f(6) = \frac{6^2-4(6)+36}{2}\] \[\Large f(6) = \frac{36-4(6)+36}{2}\] \[\Large f(6) = \frac{36-24+36}{2}\] \[\Large f(6) = \frac{12+36}{2}\] \[\Large f(6) = \frac{48}{2}\] \[\Large f(6) = 24\]

jimthompson5910 (jim_thompson5910):

So this means it costs $24 to make 6 widgets

OpenStudy (anonymous):

ohh okay so i always going to subsitute the x for its value

OpenStudy (anonymous):

and how do you get that list of numbers? like this one that you type [2,16], [4,18], [10,48] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Given: "The company also discovered that it cost $16 to produce 2 widgets, $18 to produce 4 widgets, and $48 to produce 10 widgets." "it cost $16 to produce 2 widgets" means x = 2 and y = 16 "$18 to produce 4 widgets" means x = 4 and y = 18 "$48 to produce 10 widgets" means x = 10 and y = 48

jimthompson5910 (jim_thompson5910):

all that means we know (2,16), (4, 18) and (10,48) all lie on this parabola

OpenStudy (anonymous):

and how you know which ones where on the same parentesis

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

some of the numbers where in the same parenthesis

OpenStudy (anonymous):

oh nevermind i get it. thanks for your help! i have to go Thanks a lot!!

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

alright I'm glad it's making sense now